A clock pendulum has a period of 2.00s. A simple pendulum setup in front of it gains on the clock so

      that the two vibrate in phase at intervals of 22.0s. Calculate:

      (a.) the period of the simple pendulum

      (b.) the fractional change in length of the simple pendulum necessary for the two periods to be

      equal. 



1
Expert's answer
2021-03-02T07:41:09-0500

Answer

A clock pendulum has a time period of "T_1=" 2.00s

the period of the simple pendulum

"T_2=\frac{22}{2}=11sec."

Now

the fractional change in length of the simple pendulum necessary for the two periods to be equal. 

"T=2\pi\sqrt{ \frac{l}{g}}"

Both periods are equal then

"T_1=T_2"

"2\pi\sqrt{ \frac{l_1}{g}}=2\pi\sqrt{ \frac{l_2}{g}}"

So fractional change in length

"\frac{l_2}{l_1}=(\frac{T_2}{T_1}) ^2=(\frac{11}{2}) ^2=30.25"